OpenStudy (anonymous):

-13/18divided by 20/27 perform the indicated operation

OpenStudy (anonymous):

\[-\frac{13}{18}/ \frac{20}{27}=-\frac{13}{18}*\frac{27}{20}=-\frac{39}{40}\]